### v2

# Atom-Induced Field Squeezing Predicted by Magnus Expanding the Jaynes-Cummings Model for a Two-Level Atom

Abstract
By performing up to the second order Magnus expansion to the Jaynes-Cummings Model Hamiltonian governing the interaction of a two-level atom and a single cavity mode electromagnetic field, we find the term involving the squeezing of light beyond the Rotating Wave Approximation. Our results demonstrate that the Magnus expansion provides a systematic framework to capture atom-induced squeezing effects. These findings suggest new avenues for engineering nonclassical states in simple light–matter systems without requiring multi-level structures or strong nonlinear media.
preprint: APS/123-QED
I Introduction
In general, solving the Schrodinger equation involving the Hamiltonian with an explicit time-dependence is difficult. The Magnus Expansion (ME) provides an exact solution of the time evolution operator whose logarithm is expressed as a series of time-ordered integrals of nested commutators [1, 2].
To solve the time-dependent Hamiltonian governing the light-matter interaction, one usually adapts the Rotating Wave Approximation (RWA) by ignoring the contributions of high-frequency parts of either the rotated Hamiltonian or the interaction Hamiltonian [3, 4]. There are also time-averaging approaches where one finds an approximate solution for long timescales by applying a time-averaging operator on the Hamiltonian [5, 6, 7]. Such time-averaging approaches are useful for open system quantum dynamics where the quantum system is allowed to interact with complex degrees of freedom of the environment [8]. There also exists a method where one coarse-grains the Magnus evolution operator into time steps [9].
When it comes to solving the time-dependent Hamiltonian equation governing light-matter interaction between a quantized field and a two-level atom, one first rotates the time-dependent Hamiltonian into a time-independent form, using a unitary transformation that represents a rotation [10, 11]. One well-known model of quatum light-matter interaction is the Jaynes-Cummings (JC) model, which assumes that only one cavity mode of the quantum electromagnetic field interacts with the atom via weak coupling [12, 13, 14]. In this paper, we perform up to second-order Magnus Expansion on the time-evolution operator for the JC model and make comparisons to the RWA and time-averaging methods. Moreover, we demonstrate “beyond the RWA” contributions, such as the Bloch-Siegert shift and the atom-induced squeezing of light.
The JC model has been a fundamental framework for quantum light-matter interaction. Phenomena beyond the RWA, such as the Bloch Siegert (BS) shift and the atom-induced field squeezing, have been studied using the JC model in the past. Various authors have analytically and numerically demonstrated field squeezing within the JC model framework [15, 16]. The BS shift, originally predicted by Bloch and Sigert in 1940 [17], has been experimentally demonstrated using circuit QED systems [18].
In this work, we analytically study the Jaynes–Cummings model without the rotating wave approximation using the Magnus expansion up to second order. While the Bloch–Siegert shift is a known feature of beyond-RWA dynamics, we find that the second-order expansion also predicts a squeezing-like interaction induced solely by a two-level atom. To the best of our knowledge, such an effect has not been derived in closed form from the JC model without additional nonlinearities or multi-level structure. This result demonstrates that field squeezing can emerge naturally from time-dependent perturbative dynamics, and offers a new analytic perspective on light–matter interactions beyond the RWA.
II Methods
II.1 Transforming the Hamiltonian into the Rotating Frame
[…]
Acting $\hat{U}^{\dagger}(t)$ on both sides of the equation and using the unitary condition, one gets the following new Schrodinger equation.
[…]
Defining the “rotated” Hamiltonian as in equation 4, one gets the Schrodinger equation in the “rotated” frame as shown in equation 5.
$\hat{H}_{rot}(t)=\hat{U}^{\dagger}(t)\hat{H}(t)\hat{U}(t)-i\hbar\hat{U}^{\dagger}(t)\frac{d\hat{U}(t)}{dt}$ (4)
[…]
The conditions for the solvability of general multi-level atomic systems interacting with a classical electromagnetic field under the RWA were determined in 1975 by Einwohner, Wong, and Garrison [19]. It has also been shown in our previous work that only a few of the higher-level systems are time-independent for solution via matrix exponentiation [20]. Such conditions apply to the JC model with general multi-level atomic systems, as rotating the field contributes in eliminating the field term and inducing frequency-dependent terms, as we will see in section II.3. The analysis for the JC model for general multilevel atomic systems is beyond the scope of this research.
II.2 The Magnus Expansion
In general, solving a system with a time-dependent Hamiltonian is a challenging task. One approach is to rotate the Hamiltonian into a frame where the Hamiltonian is time-independent and solve the Hamiltonian under the RWA. Even under the RWA, only a few systems can be rotated into a time-independent form [20]. However, there exists a method that gives exact solutions to any time-dependent Schrodinger equation involving a Hamiltonian with explicit time-dependence on the condition that the expansion converges [21]. Such an expansion is called the Magnus Expansion.
The Magnus Expansion was first introduced by Wilhelm Magnus in 1954 [1]. The method provides an expansion of the unitary time-evolution operator ($\hat{U}(t)=exp(\hat{\Omega}(t))$) in the form of a nested commutator of the Hamiltonian at different times. In the form $\hat{\Omega}(t)=\hat{\Omega}_{1}(t)+\hat{\Omega}_{2}(t)+$..., the first and second order representations of $\hat{\Omega}(t)$ is given as follows [2].
$\hat{\Omega}_{1}(t)=-\frac{i}{\hbar}\int_{0}^{t}\hat{H}(t_{1})dt_{1}$ (6)
$\hat{\Omega}_{2}(t)=\frac{1}{2}(-\frac{i}{\hbar})^{2}\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}[\hat{H}(t_{1}),\hat{H}(t_{2})]$ (7)
There has been work where one goes beyond the second order in the Magnus Expansion [22, 23]. For the scope of this research, the second-order is adequate to show beyond the RWA phenomenon.
II.3 The Jaynes-Cummings Model in the Rotating Frame
The Hamiltonian described by the JC model for a two-level atom can be broken down into the field part $\hat{H}_{F}(t)$, the atom part $\hat{H}_{at}(t)$, and the interaction part $\hat{H}_{I}(t)$ as follows.
$\hat{H}(t)$ $=\hat{H_{F}}(t)+\hat{H}_{at}(t)+\hat{H}_{I}(t)$ (8)
$=\hbar\omega\hat{a}^{\dagger}\hat{a}+\frac{\hbar\omega_{0}}{2}\hat{\sigma}_{z}+\hbar g(i\hat{a}^{\dagger}\hat{\sigma}_{+}+i\hat{a}^{\dagger}\hat{\sigma}_{-}+h.a.)$ (9)
In the above equation, $\hat{a}^{\dagger}$ and $\hat{a}$ represents the creation and annihilation operators of the cavity mode of interest, $\omega$ and $\omega_{0}$ the frequency of the mode and that of the atom, $\hat{\sigma}_{z}$ the Pauli-z operator for a two-level atom, $\hat{\sigma}_{\pm}$ the atomic raising and lowering operators, and g the coupling constant between the electromagnetic field and the atom transition. The matrix representation of the $\hat{\sigma}$ operators is as follows.
$\hat{\sigma}_{z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},\hat{\sigma}_{-}=\begin{pmatrix}0&0\\ 1&0\end{pmatrix},\hat{\sigma}_{+}=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}$ (10)
In standard calculations of the JC model, we perform two rotations: one for the field and one for the atom. We first rotate the atom part of the Hamiltonian by using the unitary matrix $\hat{U}(t)=\exp(-i\omega_{0}\hat{\sigma}_{z}t/2)$. Notice that $i\hbar\hat{U}^{\dagger}(t)\frac{d\hat{U}(t)}{dt}=-\hbar\omega_{0}/2\hat{\sigma}_{z}$ cancels with $\hat{H}_{at}(t)$. The Pauli-z operator $\hat{\sigma}_{z}$ commutes with the creation, annihilation operators, and itself. Using the unitary condition, we get $\hat{U}^{\dagger}(t)(\hat{H}_{F}(t)+\hat{H}_{at}(t))\hat{U}(t)=\hat{H}_{F}(t)+\hat{H}_{at}(t)$. The interaction part of the Hamiltonian gets rotated as $\hat{U}^{\dagger}(t)\hat{H}_{I}(t)\hat{U}(t)=\hbar g(\hat{a}^{\dagger}-\hat{a})\hat{\sigma}_{-}e^{-i\omega_{0}t}+(\hat{a}^{\dagger}-\hat{a})\hat{\sigma}_{+}e^{i\omega_{0}t}$. Hence, the Hamiltonian after the atomic rotation is determined as follows.
$\hat{H}_{rot}(t)=\hbar\omega\hat{a}^{\dagger}\hat{a}+i\hbar g(\hat{a}^{\dagger}-\hat{a})(\hat{\sigma}_{-}e^{-i\omega_{0}t}+\hat{\sigma}_{+}e^{i\omega_{0}t})$ (11)
Next, we do a second rotation by introducing the field-rotation unitary $\hat{V}(t)=\exp(-i\omega\hat{n}t)$ where $\hat{n}=\hat{a}^{\dagger}\hat{a}$ is the hermitian photon number operator. Due to the hermiticity, $\hat{V}^{\dagger}(t)=\exp(i\omega\hat{n}t)$. We get $i\hbar\hat{V}^{\dagger}(t)\frac{d\hat{V}(t)}{dt}=-\hbar\omega\hat{n}$ and hence, the first term in $\hat{H}_{rot}(t)$ is cancelled. Using the Baker-Campbell-Haussdorff (BCH) lemma (see appendix A), we get the relation $\hat{V}^{\dagger}(t)\hat{a}\hat{V}(t)=\hat{a}e^{-i\omega t}$ and its conjugate $\hat{V}^{\dagger}(t)\hat{a}^{\dagger}\hat{V}(t)=\hat{a}^{\dagger}e^{i\omega t}$. Combining everything, we get the fully-rotated Hamiltonian as follows.
$\begin{split}\hat{H}^{\prime}(t)=\hbar g(&i\hat{a}^{\dagger}\hat{\sigma}_{-}e^{i\Delta t}-i\hat{a}\hat{\sigma}_{+}e^{-i\Delta t}+\\ &i\hat{a}^{\dagger}\hat{\sigma}_{+}e^{i\Sigma t}-i\hat{a}\hat{\sigma}_{-}e^{-i\Sigma t})\end{split}$ (12)
Here, we defined detuning to be $\Delta=\omega-\omega_{0}$ and the sum-frequency $\Sigma=\omega+\omega_{0}$. Notice that the “rotated” Hamiltonian is similar to the interaction picture Hamiltonian. However, we are staying in the full Schrödinger picture, as the operators hold no time dependence.
Notice the terms oscillating at frequency $\Sigma$. Such terms do not contribute much to system dynamics. Physically, the term oscillating at $\Sigma$ destroys a photon and lowers the atom from excited to ground state, while the term oscillating at $\Sigma$ creates a photon and excites the atom. These processes do not obey the conservation of energy. Nevertheless, under short timescales, the energy-time uncertainty relations allow rapid fluctuations in system dynamics. Many theoretical derivations and experimental applications have neglected such highly oscillating terms by making a so-called “Rotating Wave Approximation (RWA)” [24]. Moreover, it has also been shown that one must be careful when using the RWA as the approximation can produce non-causal behaviors [25, 26, 27, 28]. There also exists a paper by Bugarth and colleagues on the effects of the RWA and the experimental validity of the approximation [29]. In this work, we briefly go over the “averaging” process of the first-order Magnus expansion and its relation to the RWA. We also capture non-trivial contributions from the terms neglected by the RWA in section III.2.
III Magnus Expansion on the Rotated Hamiltonian
III.1 The First Order Magnus Expansion
Recall that the first-order Magnus Expansion for the time-evolution operator is as follows.
$\hat{\Omega}_{1}(t)=-\frac{i}{\hbar}\int_{0}^{t}\hat{H}(t_{1})dt_{1}$ (13)
Substituting equation 12 and integrating over time, we get the first-order time evolution operator as follows.
$\begin{split}\hat{\Omega}_{1}(t)=&\frac{ig}{\Delta}(\hat{a}^{\dagger}\hat{\sigma}_{-}(1-e^{i\Delta t})+\hat{a}\hat{\sigma}_{+}(1-e^{-i\Delta t}))+\\ &\frac{ig}{\Sigma}(\hat{a}^{\dagger}\hat{\sigma}_{+}(1-e^{i\Sigma t})+\hat{a}\hat{\sigma}_{-}(1-e^{-i\Sigma t}))\end{split}$ (14)
Our solution for the time-evolution operator for the JC model has been exact so far with no approximations. However, when dealing with highly-detuned systems, $\Sigma$ is around two times the optical frequencies while $\Delta$ is small, depending on how much detuning one sets. Therefore, the $\frac{1}{\Sigma}$ terms die-off quickly while the $\frac{1}{\Delta}$ dependent terms dominates the interaction. One omits the highly oscillating terms in the Hamiltonian and makes the Rotating Wave Approximation (RWA) when dealing with near-resonance transitions. In literature, the first-order Magnus Expansion is commonly described as the RWA, and higher orders are associated with “beyond the RWA” [30]. In the next section, we will see how the second-order Magnus Expansion gives us “beyond the RWA” contributions such as the field-squeezing operator.
III.2 The Second Order Magnus Expansion
Recall that the second-order Magnus Expansion term is given as follows.
$\hat{\Omega}_{2}(t)=-\frac{1}{2\hbar^{2}}\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}[\hat{H}^{\prime}(t_{1}),\hat{H}^{\prime}(t_{2})]$ (15)
We have to find the commutator at different times. First, we write down the well-known commutation relations.
$\begin{split}&[\hat{a},\hat{a}^{\dagger}]=1,\\ &[\hat{\sigma}_{-},\hat{\sigma}_{+}]=-\hat{\sigma}_{z}\end{split}$ (16)
Noting that every commutator commutes with itself, we get the commutator at two different times as follows.
$\begin{split}-\frac{[\hat{H}^{\prime}(t_{1}),\hat{H}^{\prime}(t_{2})]}{\hbar^{2}g^{2}}=&[\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}\hat{\sigma}_{+}](-e^{i\Delta(t_{1}-t_{2})}+e^{-i\Delta(t_{1}-t_{2})})+\\ &[\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}^{\dagger}\hat{\sigma}_{+}](e^{i(\Delta t_{1}+\Sigma t_{2})}-e^{i(\Sigma t_{1}+\Delta t_{2})})+\\ &[\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}\hat{\sigma}_{-}](-e^{i(\Delta t_{1}-\Sigma t_{2})}+e^{-i(\Sigma t_{1}-\Delta t_{2})})+\\ &[\hat{a}^{\dagger}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{+}](-e^{i(\Sigma t_{1}-\Delta t_{2})}+e^{-i(\Delta t_{1}-\Sigma t_{2})})+\\ &[\hat{a}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{-}](e^{-i(\Delta t_{1}+\Sigma t_{2})}-e^{-i(\Sigma t_{1}+\Delta t_{2})})+\\ &[\hat{a}^{\dagger}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{-}](-e^{i\Sigma(t_{1}-t_{2})}+e^{-i\Sigma(t_{1}-t_{2})})\end{split}$ (17)
Since the commutators do not have explicit time dependence, we do the time integrals first and see the contributions of each term present in equation 17. The time integrals are as follows.
$I_{1}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(-e^{i\Delta(t_{1}-t_{2})}+e^{-i\Delta(t_{1}-t_{2})})$
$=-\frac{2it}{\Delta}+\frac{2i}{\Delta^{2}}sin(\Delta t)$ (18)
$I_{2}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(e^{i(\Delta t_{1}+\Sigma t_{2})}-e^{i(\Sigma t_{1}+\Delta t_{2})})$
$=\frac{\omega_{0}e^{2i\omega t}-\omega e^{i\Sigma t}+\omega e^{i\Delta t}-\omega_{0}}{\omega(\omega^{2}-\omega_{0}^{2})}$ (19)
$I_{3}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(-e^{i(\Delta t_{1}-\Sigma t_{2})}+e^{-i(\Sigma t_{1}-\Delta t_{2})})$ (20)
$I_{4}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(-e^{i(\Sigma t_{1}-\Delta t_{2})}+e^{-i(\Delta t_{1}-\Sigma t_{2})})$ (21)
$I_{5}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(e^{-i(\Delta t_{1}+\Sigma t_{2})}-e^{-i(\Sigma t_{1}+\Delta t_{2})})$
$=\frac{\omega_{0}e^{-2i\omega t}-\omega e^{-i\Sigma t}+\omega e^{-i\Delta t}-2\omega_{0}}{(\omega^{2}-\omega_{0}^{2})\omega}=I_{2}^{*}$ (22)
$I_{6}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(-e^{i\Sigma(t_{1}-t_{2})}+e^{-i\Sigma(t_{1}-t_{2})})$
$=-\frac{2it}{\Sigma}+\frac{2i}{\Sigma^{2}}sin(\Sigma t)$ (23)
Note that $I_{3}$ and $I_{4}$ are not evaluated because the commutators associated with these integrals are zero. Also, $I_{2}$ and $I_{5}$ being complex conjugates of each other is crucial in yielding the atomic squeezing term later. For notation, call $I_{2}=\zeta$ and $I_{5}=\zeta^{*}$.
A useful identity for the commutator of products is as follows.
$\hat{A}\hat{B},\hat{C}\hat{D}]=\hat{A}[\hat{B},\hat{C}]\hat{D}+\hat{A}\hat{C}[\hat{B},\hat{D}]+[\hat{A},\hat{C}]\hat{D}\hat{B}+\hat{C}[\hat{A},\hat{D}]\hat{B}$ (24)
Using the commutator identity, we get the following commutator relations.
$|\displaystyle[\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}\hat{\sigma}_{+}]=-\hat{a}^{\dagger}\hat{a}\hat{\sigma}_{z}-\hat{\sigma}_{+}\hat{\sigma}_{-}=-\hat{a}^{\dagger}\hat{a}\hat{\sigma}_{z}-\ket{1}\bra{1}$ (25)
$[\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}^{\dagger}\hat{\sigma}_{+}]=-(\hat{a}^{\dagger})^{2}\hat{\sigma}_{z}$ (26)
$[\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}\hat{\sigma}_{-}]=-\hat{\sigma}_{-}^{2}=0$ (27)
$[\hat{a}^{\dagger}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{+}]=-\hat{\sigma}_{+}^{2}=0$ (28)
$[\hat{a}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{-}]=\hat{a}^{2}\hat{\sigma}_{z}$ (29)
$|\displaystyle[\hat{a}^{\dagger}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{-}]=\hat{a}^{\dagger}\hat{a}\hat{\sigma}_{z}-\hat{\sigma}_{-}\hat{\sigma}_{+}=\hat{a}^{\dagger}\hat{a}\hat{\sigma}_{z}-\ket{0}\bra{0}$ (30)
[…]
Notice that the second-order Magnus operator is manifestly anti-hermitian as the first two terms involve the product of a purely imaginary number and a hermitian operator, while the last term is the well known anti-hermitian squeezing operator. This agrees with the condition that every order of $\hat{\Omega}(t)$ must be anti-hermitian to maintain unitarity of the Magnus time evolution operator. The term $i\frac{g^{2}t}{\Delta}\hat{a}^{\dagger}\hat{a}\hat{\sigma}_{z}$ represents a process where the atomic energy is shifted according to the number of photons present. This can be interpreted as an AC-Stark Shift. Since $|\hat{\sigma}_{+}\hat{\sigma}_{-}=\ket{1}\bra{1}$, it represents the frequency shift of the excited state. The term $i\frac{g^{2}t}{\Sigma}\hat{a}^{\dagger}\hat{a}\hat{\sigma}_{z}$ represents the Bloch-Siegert shift as it is a small photon-induced energy shift suppressed by a large frequency $\omega+\omega_{0}$. Moreover, our results agree with Wall’s prediction that the Bloch-Siegert shift scales with the number of photons, which was later verified experimentally [31, 32].
The last term involves the $\zeta^{*}\hat{a}^{2}-\zeta(\hat{a}^{\dagger})^{2}$ term, which resembles the squeezing operator of the quantized electromagnetic field with $\zeta=\frac{\omega_{0}e^{2i\omega t}-\omega e^{i\Sigma t}+\omega e^{i\Delta t}-\omega_{0}}{\omega(\omega^{2}-\omega_{0}^{2})}$ [14]. Therefore, it can be interpreted as the presence of the atom induces some squeezing towards the existing light field. This behavior would not be captured with RWA, as the cross-contribution of rotating and counterrotating terms is responsible for squeezing. Moreover, the amplitude of the squeezing term is not negligible when the laser is tuned near resonance, and it seems like it diverges on resonance. However, the term does not diverge on resonance, and the proof is shown in appendix B. Inspecting the limit shown in appendix B, it can also be noted that the amplitude is quite small when the system is tuned near resonance. One interesting remark is that the coefficient of the squeezing term is an exponent that is on another exponent due to the nature of the Magnus evolution operator. This means that the squeezing effect fluctuates very rapidly in time. Therefore, it is interesting to inspect the contribution of such a term in atom dynamics for potential experimental evidence.
Notice the increase in powers of the coupling constant in the Magnus expansion. The Magnus expansion converges when $gt<<\pi$ [2]. The standard values for coupling constants for cavity QED setups are $g=10^{6}-10^{7}$Hz [29]. For long timescale quantum dynamics, coarse-graining the Magnus expansion on periods on the orders below the microsecond scale has been suggested [9].
IV Conclusion
In this paper, we kept all the terms of the Jaynes-Cummings Hamiltonian by not performing the RWA prematurely. We performed first and second order Magnus Expansion for the time-evolution operator using the Jaynes-Cummings Hamiltonian rotated in both atom and field frames. We have seen why the RWA is valid from the first-order Magnus expansion. On the other hand, we have seen the non-trivial contributions from the counterrotating terms, such as the Bloch-Siegert Shift and the creation of squeezed light due to the presence of atoms. The key contribution of this paper is the analytic prediction of field squeezing from a two-level atom in the JC model using the Magnus expansion. This term, emerging beyond the RWA, offers a new theoretical insight into cavity QED systems.
[…]
Our long-term goal is to physically explain the nature of RWA by finding its deviations from the exact Magnus Expansion approach. This work highlights how the Magnus Expansion not only justifies the use of the RWA but also reveals non-negligible quantum effects, offering a physical framework to understand the effects beyond the RWA.
Acknowledgements.
I would like to express my appreciation to Professor Daniel F. V. James from the Department of Physics, University of Toronto, for reading the manuscript and providing helpful discussions and comments.
Disclosure of Interest
The author reports that there are no competing interests to declare.
Appendix A BCH identities
For arbitrary operators $\hat{A},\hat{B}$ and an arbitrary complex number $\zeta$, the Baker-Campbell-Hausdorff (BCH) lemma states the following identity.
$exp(\zeta\hat{A})\ \hat{B}\ exp(-\zeta\hat{A})=\hat{B}+\zeta[\hat{A},\hat{B}]+\frac{\zeta^{2}}{2!}[\hat{A},[\hat{A},\hat{B}$]]+...
Using the BCH lemma, we derive the following identities.
$\hat{V}^{\dagger}\hat{a}^{\dagger}\hat{V}$ $=exp(i\omega\hat{a}^{\dagger}\hat{a}t)\hat{a}^{\dagger}exp(-i\omega\hat{a}^{\dagger}\hat{a}t)$
$=\hat{a}^{\dagger}+i\omega t[\hat{a}^{\dagger}\hat{a},\hat{a}^{\dagger}]+\frac{(i\omega t)^{2}}{2}[\hat{a}^{\dagger}\hat{a},[\hat{a}^{\dagger}\hat{a},\hat{a}^{\dagger}]]$
Since $\hat{a}^{\dagger}\hat{a},\hat{a}^{\dagger}]=\hat{a}^{\dagger}$, we get the following series.
[…]
Doing the same for the annihilation operator, we get the following identity.
[…]
Using the Taylor series expressions for exponents, we get the necessary expressions to perform the light-rotations as follows.
$\hat{V}^{\dagger}\hat{a}^{\dagger}\hat{V}=\hat{a}^{\dagger}e^{i\omega t}$ (32)
$\hat{V}^{\dagger}\hat{a}\hat{V}=\hat{a}e^{-i\omega t}$ (33)
Appendix B Proof that the Coefficient of the Squeezing Term does not Diverge on Resonance
Recall that the coefficient of the squeezing term is as follows.
$\zeta=\frac{g^{2}}{2}\frac{\omega_{0}e^{2i\omega t}-\omega e^{i\Sigma t}+\omega e^{i\Delta t}-\omega_{0}}{(\omega^{2}-\omega_{0}^{2})\omega}$ (34)
On resonance, $\omega=\omega_{0}$. This means $\Sigma=\omega+\omega_{0}=2\omega$ and $\Delta=\omega-\omega_{0}=0$. On the limit of resonance, the constant can be represented as follows.
$\zeta$ $=\frac{g^{2}}{2}\frac{(\omega_{0}-\omega)e^{2i\omega t}+(\omega-\omega_{0})}{\omega(\omega-\omega_{0})(\omega+\omega_{0})}$
$=\frac{g^{2}}{2}\frac{(\omega-\omega_{0})(1-e^{2i\omega t})}{\omega(\omega-\omega_{0})(\omega+\omega_{0})}$ (35)
On a limit that $\omega$ approaches $\omega_{0}$, the terms $\omega-\omega_{0}$ cancel and the constant reduces to a non-divergent form on resonance.
$\lim_{\omega\to\omega_{0}}\zeta=\frac{g^{2}}{2}\frac{1-e^{2i\omega t}}{\omega(\omega+\omega_{0})}$ (36)
Since $\zeta$ does not diverge on resonance, its complex conjugate $\zeta^{*}$ does not diverge either.