### v4

# Conditional squeezing induced by a two-level system: a first-principles approach to QND readout

Abstract
We present a systematic Magnus expansion treatment of light-matter interaction beyond the Rotating Wave Approximation. We show that at the second order of Magnus series, the time-evolution operator acquires both energy-shifts and squeezing contributions. In addition to the energy shifts caused by vacuum and photon numbers, the second-order evolution operator contains a term that induces conditional squeezing of the field mode depending on the state of the atom. Such a term suggests a natural mechanism for phase-sensitive, quantum non-demolishing type measurements of squeezed light via homodyne detection. We also show that the second-order Magnus operator in a close SU(1,1) algebra, ensuring the exponentiation of the Magnus series yields a well-defined unitary evolution. By deriving squeezing directly from the Jaynes-Cummings Hamiltonian, our results clarify how energy shifts and $\hat{\sigma}_{z}$-dependent squeezing arise from an SU(1,1) structure, with direct implications for phase-sensitive and quantum nondemolition measurements.
Magnus expansion, non-demolishing measurements, quadrature squeezing, Jaynes-Cummings model, beyond the Rotating Wave Approximation (RWA), SU(1,1) group
preprint: APS/123-QED
I Introduction
Squeezed states of light, first predicted theoretically to reduce quantum fluctuations beyond the shot-noise limit [6], were later experimentally realized by four-wave mixing in an optical cavity [23]. Such states were later incorporated in enhanced precision measurements in interferometric setups [28, 25, 22]. Beyond laboratory-scale demonstrations, squeezed states have also been applied in large-scale interferometric detectors. The LIGO collaboration has demonstrated the use of squeezed light to enhance the sensitivity of gravitational detectors [25].
Most measurements in quantum information destroy the state of the qubit due to the collapse postulate. It is important to have a way of performing measurements from a quantum system without irreversibly disturbing the observable of interest. Such measurements are called quantum nondemolishing (QND) measurements [19, 4]. In this article, we demonstrate how the interaction of a simple two-level system with a single-mode field opens possibilities of performing such measurements.
We perform up to second-order Magnus Expansion on the time-evolution operator for the Jaynes-Cummings(JC) model [11]. We demonstrate “beyond the Rotating Wave Approximation(RWA)” contributions, such as the Bloch-Siegert shift and the atom-induced conditional squeezing of light. The JC model has been a fundamental framework for quantum light-matter interaction. Various authors have analytically and numerically demonstrated field squeezing within the JC model framework [12, 30]. The Bloch-Siegert shift, originally predicted by Bloch and Siegert in 1940 [3], has been experimentally demonstrated using circuit QED systems [8]. While the Bloch–Siegert shift is a known feature of beyond-RWA dynamics, we find that the second-order expansion also predicts a two-photon term governing the squeezing of light induced solely by a two-level atom.
Although squeezing effects in two-level systems have been discussed in the literature, our approach derives the effect in closed form from first principles, without recourse to multilevel extensions or external nonlinearities. Traditionally, squeezed light is generated in nonlinear optical media using techniques such as degenerate parametric down-conversion in $\chi^{(2)}$ crystals in an optical cavity [29, 31] and four-wave mixing [24, 21]. By contrast, we demonstrate that even the simplest Jaynes–Cummings two-level atom intrinsically contains a squeezing term when treated beyond the RWA. While the magnitude of the squeezing is not expected to compete with engineered nonlinear media, the finding highlights an overlooked resource. In particular, deriving a squeezing coefficient directly from first principles strengthens the value of theoretical quantum optics, where simplistic models not only predict nonclassical light but also provide a foundation for understanding how such effects can be systematically extended to more complex, scalable systems.
This paper is arranged as follows. In Section II, we present a brief discussion of textbook materials such as the Magnus Expansion and the Jaynes-Cummings Model. In Section III, we present our calculation of the second-order Magnus operator and its implications toward QND Measurement by homodyne detection of squeezed light. Finally, Section IV summarizes our conclusions and highlights future work.
II Methods
II.1 The Magnus Expansion
The Magnus Expansion was first introduced by Wilhelm Magnus in 1954 [14]. The method provides an expansion of the unitary time-evolution operator ($\hat{U}(t)=exp(\hat{\Omega}(t))$) in the form of a nested commutator of the Hamiltonian at different times. In the form $\hat{\Omega}(t)=\hat{\Omega}_{1}(t)+\hat{\Omega}_{2}(t)+$..., the first and second order representations of $\hat{\Omega}(t)$ is given as follows [2].
$\hat{\Omega}_{1}(t)=-\frac{i}{\hbar}\int_{0}^{t}\hat{H}(t_{1})dt_{1}$ (1)
$\hat{\Omega}_{2}(t)=\frac{1}{2}(-\frac{i}{\hbar})^{2}\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}[\hat{H}(t_{1}),\hat{H}(t_{2})]$ (2)
There has been work where one goes beyond the second order in the Magnus Expansion [1, 13]. For the scope of this research, the second-order is adequate to show beyond the RWA phenomenon.
II.2 The Jaynes-Cummings Model in the Rotating Frame
The Hamiltonian described by the JC model for a two-level atom can be broken down into the field part $\hat{H}_{F}(t)$, the atom part $\hat{H}_{at}(t)$, and the interaction part $\hat{H}_{I}(t)$ as follows.
$\hat{H}(t)$ $=\hat{H_{F}}(t)+\hat{H}_{at}(t)+\hat{H}_{I}(t)$ (3)
$=\hbar\omega\hat{a}^{\dagger}\hat{a}+\frac{\hbar\omega_{0}}{2}\hat{\sigma}_{z}+\hbar g(i\hat{a}^{\dagger}\hat{\sigma}_{+}+i\hat{a}^{\dagger}\hat{\sigma}_{-}+h.a.)$ (4)
Here $\hbar$ represents the Planck’s constant divided by $2\pi$, i the imaginary number, g the coupling strength between the atom and the field, $\omega$ the frequency of the cavity field mode, $\omega_{0}$ the transition frequency of the two-level system, $\hat{a}^{\dagger}$ and $\hat{a}$ the photon creation and annihilation operators f the field, $\hat{\sigma}_{z}$ the Pauli-z operator of the atom, and $\hat{\sigma}_{+}$ and $\hat{\sigma}_{-}$ the raising and lowering operator in the atomic Hilbert space. The term $h.a.$ represents the Hermitian adjoint of the operator in brackets. Defining the “rotated” Hamiltonian as in equation 5, one can write down the Schrodinger equation in the “rotated frame” [9].
$\hat{H}_{rot}(t)=\hat{U}^{\dagger}(t)\hat{H}(t)\hat{U}(t)-i\hbar\hat{U}^{\dagger}(t)\frac{d\hat{U}(t)}{dt}$ (5)
Note that we stay in Schrodinger’s picture in this article as only the state gains time-dependence via a unitary rotation. The operators obtain no time dependence in contrast to the interaction picutre.
In standard calculations of the JC model, we perform two rotations: one for the field and one for the atom [9]. The unitary operator representing the atomic rotation is $\hat{U}(t)=\exp(-i\omega_{0}\hat{\sigma}_{z}t/2)$ and that of the field is $\hat{V}(t)=\exp(i\omega\hat{n}t)$ where $\hat{n}$ is the hermitian photon number operator defined as $\hat{n}=\hat{a}^{\dagger}\hat{a}$. After the Hamiltonian rotations using equation 5, we get the fully rotated Jaynes-Cummings Hamiltonian as follows.
$\begin{split}\hat{H}^{\prime}(t)=\hbar g(&i\hat{a}^{\dagger}\hat{\sigma}_{-}e^{i\Delta t}-i\hat{a}\hat{\sigma}_{+}e^{-i\Delta t}+\\ &i\hat{a}^{\dagger}\hat{\sigma}_{+}e^{i\Sigma t}-i\hat{a}\hat{\sigma}_{-}e^{-i\Sigma t})\end{split}$ (6)
Here, we defined detuning to be $\Delta=\omega-\omega_{0}$ and the sum-frequency $\Sigma=\omega+\omega_{0}$. Notice the terms oscillating at frequency $\Sigma$. Such terms do not contribute much to system dynamics. Physically, the term oscillating at $\Sigma$ destroys a photon and lowers the atom from excited to ground state, while the term oscillating at $\Sigma$ creates a photon and excites the atom. These processes do not obey the conservation of energy. Nevertheless, under short timescales, the energy-time uncertainty relations allow rapid fluctuations in system dynamics. Many theoretical derivations and experimental applications have neglected such highly oscillating terms by making a so-called “Rotating Wave Approximation (RWA)” [18]. Time-averaging effective Hamiltonian theories have also made similar approximations to the RWA by neglecting higher-order fluctuations by taking ensemble averages of the dynamics [10]. Moreover, it has also been shown that one must be careful when using the RWA as the approximation can produce non-causal behaviors [16, 15, 20, 7]. There also exists a paper by Bugarth and colleagues on the effects of the RWA and the experimental validity of the approximation [5]. In this work, we briefly go over the “averaging” process of the first-order Magnus expansion and its relation to the RWA. We also capture non-trivial contributions from the terms neglected by the RWA in section III.2.
III Magnus Expansion on the Rotated Hamiltonian
III.1 The First Order Magnus Expansion
Recall that the first-order Magnus Expansion for the time-evolution operator is as follows.
$\hat{\Omega}_{1}(t)=-\frac{i}{\hbar}\int_{0}^{t}\hat{H}(t_{1})dt_{1}$ (7)
Substituting equation 6 and integrating over time, we get the first-order time evolution operator as follows.
$\begin{split}\hat{\Omega}_{1}(t)=&\frac{ig}{\Delta}(\hat{a}^{\dagger}\hat{\sigma}_{-}(1-e^{i\Delta t})+\hat{a}\hat{\sigma}_{+}(1-e^{-i\Delta t}))+\\ &\frac{ig}{\Sigma}(\hat{a}^{\dagger}\hat{\sigma}_{+}(1-e^{i\Sigma t})+\hat{a}\hat{\sigma}_{-}(1-e^{-i\Sigma t}))\end{split}$ (8)
Our solution for the time-evolution operator for the JC model has been exact so far with no approximations. However, when dealing with highly-detuned systems, $\Sigma$ is around two times the optical frequencies while $\Delta$ is small, depending on how much detuning one sets. Therefore, the $\frac{1}{\Sigma}$ terms die-off quickly while the $\frac{1}{\Delta}$ dependent terms dominates the interaction. One omits the highly oscillating terms in the Hamiltonian and makes the Rotating Wave Approximation (RWA) when dealing with near-resonance transitions. In literature, the first-order Magnus Expansion is commonly described as the RWA, and higher orders are associated with “beyond the RWA” [27]. In the next section, we will see how the second-order Magnus Expansion gives us “beyond the RWA” contributions such as the field-squeezing operator.
III.2 The Second Order Magnus Expansion
Recall that the second-order Magnus Expansion term is given as follows.
$\hat{\Omega}_{2}(t)=-\frac{1}{2\hbar^{2}}\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}[\hat{H}^{\prime}(t_{1}),\hat{H}^{\prime}(t_{2})]$ (9)
We have to find the commutator at different times. Noting that every commutator commutes with itself, we get the commutator at two different times as follows.
$\begin{split}-\frac{[\hat{H}^{\prime}(t_{1}),\hat{H}^{\prime}(t_{2})]}{\hbar^{2}g^{2}}&=[\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}\hat{\sigma}_{+}](-e^{i\Delta(t_{1}-t_{2})}+e^{-i\Delta(t_{1}-t_{2})})+\\ &[\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}^{\dagger}\hat{\sigma}_{+}](e^{i(\Delta t_{1}+\Sigma t_{2})}-e^{i(\Sigma t_{1}+\Delta t_{2})})+\\ &[\hat{a}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{-}](e^{-i(\Delta t_{1}+\Sigma t_{2})}-e^{-i(\Sigma t_{1}+\Delta t_{2})})+\\ &[\hat{a}^{\dagger}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{-}](-e^{i\Sigma(t_{1}-t_{2})}+e^{-i\Sigma(t_{1}-t_{2})})\end{split}$ (10)
Here, we neglect terms involving $\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}\hat{\sigma}_{-}$] and $\hat{a}^{\dagger}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{+}$] as the commutators evaluate to zero. Since the commutators do not have explicit time dependence, we do the time integrals first and see the contributions of each term present in equation 10. The time integrals are as follows.
$I_{1}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(-e^{i\Delta(t_{1}-t_{2})}+e^{-i\Delta(t_{1}-t_{2})})$ (11)
$=-\frac{2it}{\Delta}+\frac{2i}{\Delta^{2}}sin(\Delta t)$ (12)
$I_{2}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(e^{i(\Delta t_{1}+\Sigma t_{2})}-e^{i(\Sigma t_{1}+\Delta t_{2})})$ (13)
$=\frac{\omega_{0}e^{2i\omega t}-\omega e^{i\Sigma t}+\omega e^{i\Delta t}-\omega_{0}}{\omega(\omega^{2}-\omega_{0}^{2})}$ (14)
$I_{3}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(e^{-i(\Delta t_{1}+\Sigma t_{2})}-e^{-i(\Sigma t_{1}+\Delta t_{2})})$ (15)
$=\frac{\omega_{0}e^{-2i\omega t}-\omega e^{-i\Sigma t}+\omega e^{-i\Delta t}-\omega_{0}}{(\omega^{2}-\omega_{0}^{2})\omega}=I_{2}^{*}$ (16)
$I_{4}$ $=\int_{0}^{t}dt_{1}\int_{0}^{t_{1}}dt_{2}(-e^{i\Sigma(t_{1}-t_{2})}+e^{-i\Sigma(t_{1}-t_{2})})$ (17)
$=-\frac{2it}{\Sigma}+\frac{2i}{\Sigma^{2}}sin(\Sigma t)$ (18)
Note that $I_{2}$ and $I_{3}$ being complex conjugates of each other is crucial in yielding the atomic squeezing term later. For notation, call $I_{2}=\zeta(t)/g^{2}$ and $I_{3}=\zeta^{*}(t)/g^{2}$.
A useful identity for the commutator of products is as follows.
$\hat{A}\hat{B},\hat{C}\hat{D}]=\hat{A}[\hat{B},\hat{C}]\hat{D}+\hat{A}\hat{C}[\hat{B},\hat{D}]+[\hat{A},\hat{C}]\hat{D}\hat{B}+\hat{C}[\hat{A},\hat{D}]\hat{B}$ (19)
Using the commutator identity, we get the following commutator relations.
$|\displaystyle[\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}\hat{\sigma}_{+}]=-\hat{a}^{\dagger}\hat{a}\hat{\sigma}_{z}-\hat{\sigma}_{+}\hat{\sigma}_{-}=-\hat{a}^{\dagger}\hat{a}\hat{\sigma}_{z}-\ket{1}\bra{1}$ (20)
$[\hat{a}^{\dagger}\hat{\sigma}_{-},\hat{a}^{\dagger}\hat{\sigma}_{+}]=-(\hat{a}^{\dagger})^{2}\hat{\sigma}_{z}$ (21)
$[\hat{a}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{-}]=\hat{a}^{2}\hat{\sigma}_{z}$ (22)
$|\displaystyle[\hat{a}^{\dagger}\hat{\sigma}_{+},\hat{a}\hat{\sigma}_{-}]=\hat{a}^{\dagger}\hat{a}\hat{\sigma}_{z}-\hat{\sigma}_{-}\hat{\sigma}_{+}=\hat{a}^{\dagger}\hat{a}\hat{\sigma}_{z}-\ket{0}\bra{0}$ (23)
[…]
Notice that the second-order Magnus operator is manifestly anti-hermitian as the first two terms involve the product of a purely imaginary number and a hermitian operator, while the last term is the well known anti-hermitian squeezing operator formally defined as $\exp(\frac{1}{2}(\zeta^{*}(t)\hat{a}^{2}-\zeta(t)(\hat{a}^{\dagger})^{2})$. This agrees with the condition that every order of $\hat{\Omega}(t)$ must be anti-hermitian to maintain unitarity of the Magnus time evolution operator.
In equation 24, we can rewrite $|\ket{1}\bra{1}$ as $\frac{1}{2}(\hat{I}_{at}+\hat{\sigma}_{z})$ and $|\ket{0}\bra{0}$ as $\frac{1}{2}(\hat{I}_{at}-\hat{\sigma}_{z})$ where $\hat{I}_{at}$ is the identity operator in the Hilbert Space of the two-level atom. Noting that the identity operator commutes with every other operator, we can ignore the contribution of the term associated with the identity operator as it only adds a global phase to the system. Hence, we rewrite equation 24 as follows.
$\begin{split}\hat{\Omega}_{2}(t)=&ig^{2}(\hat{a}^{\dagger}\hat{a}+\frac{1}{2})(\frac{t}{\Delta}-\frac{t}{\Sigma}+\frac{\sin(\Sigma t)}{\Sigma^{2}}-\frac{\sin(\Delta t)}{\Delta^{2}})\hat{\sigma}_{z}\\ &+\frac{1}{2}\left(\zeta^{*}(t)\hat{a}^{2}-\zeta(t)(\hat{a}^{\dagger})^{2}\right)\hat{\sigma}_{z}\end{split}$ (25)
The first term represents the time-dependent frequency shifts, involving the AC-stark shift governed by terms involving $\Delta$, the Bloch-Siegert shift governed by terms involving $\Sigma$. The last term involves the $\frac{1}{2}(\zeta^{*}\hat{a}^{2}-\zeta(\hat{a}^{\dagger})^{2})$ term, which resembles the squeezing operator of the quantized electromagnetic field with the closed-form expression of the squeezing coefficient presented in equation 26 [9].
$\zeta(t)=g^{2}\frac{\omega_{0}e^{2i\omega t}-\omega e^{i\Sigma t}+\omega e^{i\Delta t}-\omega_{0}}{\omega(\omega^{2}-\omega_{0}^{2})}$ (26)
Therefore, it can be interpreted as the presence of the atom induces some squeezing towards the existing light field. This behavior would not be captured with RWA, as the cross-contribution of rotating and counterrotating terms is responsible for squeezing. We note that exponentiation of Equation 25 requires care due to non-commuting operators; details are given in Appendix A.
Moreover, the dependence of the $\hat{\sigma}_{z}$ operator on the squeezing term means that the squeezed light is going to have a $\pi$ phase shift depending on whether the atom is on the ground state or the excited state. There exists a possibility that one can measure the light interacting with a two-level system in a homodyne detector to possibly determine the state of the atom using quantum non-demolition (QND) measurements for quantum information [19, 4].
IV Conclusion
In this work, we kept all the terms of the Jaynes-Cummings Hamiltonian by not performing the RWA prematurely. We performed first and second order Magnus Expansion for the time-evolution operator using the Jaynes-Cummings Hamiltonian rotated in both atom and field frames. We have seen the non-trivial contributions from the counterrotating terms, such as the Bloch-Siegert Shift and the creation of squeezed light due to the presence of a two-level atom. The key contribution of this paper is the analytic prediction of field squeezing from a two-level atom in the JC model using the Magnus expansion. Such terms, emerging beyond the RWA, not only offers a new theoretical insight into engineering squeezed states of light but also opens doors for QND measurements using simple systems. Moreover, we see that the energy shifts such as the AC-Stark shift, Bloch-Siegert shift, and the vacuum energy shift are time-dependent, by a function of f(t). It is interesting to go beyond the second order to see what other non-trivial contributions can be witnessed by the non-RWA term. It is also interesting to extend this research and determine up to second-order Magnus expansion using general multilevel atomic systems under the dipole selection rules. Future research should also study the feasibility of using a two-level system for QND measurements.
The prediction of a squeezing coefficient directly from the JC Hamiltonian emphasizes the broader significance of developing theoretical quantum optics from first principles. By deriving squeezing behavior without recourse to phenomenological assumptions, this work reinforces the fundamental link between atom–field interactions and nonclassical light. While a single two-level atom cannot serve as a practical source of squeezed states, the analytical framework presented provides a clear theoretical quantification of the atom-induced squeezing coefficient. Beyond its immediate implications, this approach highlights that rigorous first-principles treatments not only sharpen our understanding of quantum optical processes but also connect to the well-studied literature on the SU(1,1) representation of field squeezing and fluctuations.
Appendix A Commutation Relations of Two-Photon Operators
One normally expects the higher-order terms in a perturbative series to have small corrections to the squeezing coefficient. However, the non-commuting structure embedded in the second order itself induces corrections to the squeezing coefficient without needing to go to higher orders. To demonstrate this, recall the Baker-Campbell-Hausdorff(BCH) theorem [26].
[…]
Notice that the commutation relations of the individual field operators in equation 25 are as follows.
$[\frac{1}{2}(\hat{a}^{\dagger}\hat{a}+\frac{1}{2}),\frac{1}{2}\hat{a}^{2}]=-\frac{1}{2}\hat{a}^{2}$ (28)
$[\frac{1}{2}(\hat{a}^{\dagger}\hat{a}+\frac{1}{2}),\frac{1}{2}(\hat{a}^{\dagger})^{2}]=\frac{1}{2}(\hat{a}^{\dagger})^{2}$ (29)
$[\frac{1}{2}(\hat{a}^{\dagger})^{2},\frac{1}{2}\hat{a}^{2}]=-(\hat{a}^{\dagger}\hat{a}+\frac{1}{2})$ (30)
When expanding the exponent using the BCH theorem, one quickly starts noticing that every commutator one evaluates, extra terms of $(\hat{a}^{\dagger})^{2},\hat{a}^{2},\hat{a}^{\dagger}\hat{a}$ are produced. The result is a never-ending loop with every term providing extra squeezing coefficients. However, there is a compact way to evaluate the exponent since the field operators form an SU(1,1) group.
Three operators $\{\hat{K}_{0},\hat{K}_{\pm}\}$ forms an SU(1,1) group if obey the following commutation relations [17].
$[\hat{K}_{0},\hat{K}_{\pm}]=\pm\hat{K}_{\pm}$ (31)
$[\hat{K}_{+},\hat{K}_{-}]=-2\hat{K}_{0}$ (32)
Defining $\hat{K_{0}}=\frac{1}{2}(\hat{a}^{\dagger}\hat{a}+\frac{1}{2})$, $\hat{K}_{+}=\frac{1}{2}(\hat{a}^{\dagger})^{2}$ and $\hat{K}_{-}=\frac{1}{2}\hat{a}^{2}$, and we can see that the operators form an SU(1,1) group. Then we can rewrite equation 25 as follows by defining $\alpha(t)=2ig^{2}f(t)\hat{\sigma}_{z}$, and $\beta(t)=-\hat{\sigma}_{z}\zeta(t)$.
[…]
It is shown in Truax’s paper that equation 33 can be disentangled into the following form[26].
$e^{\hat{\Omega}_{2}(t)}=$ $\exp\left[p_{2}\hat{K}_{+}\right]\exp[p_{0}\hat{K}_{0}]\exp[p_{1}\hat{K}_{-}$], (34)
The coefficients $p_{i}(\alpha(t),\beta(t))$ were evaluated in explicit closed forms in Truax’s article [26]. Following Truax’s discussion, we conclude that the second-order Magnus operator resolves in a compact, closed-form unitary evolution.